OpenStudy (davidusa):

MEDAL FOR WHOEVER HELPS. Sam, Pam, Ken and Jen are exchanging gifts for the holiday. Each name is written on a slip of paper and placed in a bowl. Sam is the first to randomly select a name from the bowl. How many different gift exchange pairings are there among Sam, Pam, Ken and Jen in which each person randomly selects a name other than their own name?

OpenStudy (davidusa):

@Twitchie @adrynicoleb Please help me.

OpenStudy (anonymous):

Well, Sam can pick from 3 people, Pam can pick from 3, Ken can pick from 2, Jen can pick from 1.

OpenStudy (davidusa):

Thats what I thought wio. But the answer is 9. btw this is a mathcounts question.

OpenStudy (anonymous):

Well, \(3+3+2+1 =9\), though it is a bit strange.

OpenStudy (davidusa):

wio, adrynicoleb, I have to go eat dinner at a restaurant. I'll be back in a long time so please keep answering but dont expect me to reply.

OpenStudy (davidusa):

i'll probably be back at 6:30 maybe

OpenStudy (davidusa):

wio, here is the explanation i got. We can determine the answer by making an organized list. Suppose the friends select a name from the bowl in this order S, P, K and J. The table below shows the different pairings in which each person selects a name other than their own name: There are 9 such gift exchange pairings.

OpenStudy (davidusa):

table did not show up :( ill just draw it.

OpenStudy (davidusa):

gtg now wio. ill give u a medal when i get back. thanks for helping.

OpenStudy (davidusa):

but please keep trying to answer it.

OpenStudy (wolf1728):

Here's how it is done: There are 4 factorial (4!) ways in which 4 people can choose 4 names SAM PAM KEN JEN To make things easier to track, we'll call Sam 1, Pam 2, Ken 3, and Jen 4 The attached graphic sows all 24 ways that 4 people can choose 4 names. The ones marked off with a red arrow indicate the 9 ways in which none of the people choose their own names.

OpenStudy (davidusa):

@wolf1728 come

OpenStudy (davidusa):

hi wio and honorsstudent.

OpenStudy (davidusa):

Someone please reply

OpenStudy (wolf1728):

Hello

OpenStudy (davidusa):

hi wolf can i have a quick way to do it? (this is a mathcounts problem so i need to solve it quickly)

OpenStudy (wolf1728):

You mean the gift problem resulting in 9?

OpenStudy (davidusa):

yea thats the problem

OpenStudy (davidusa):

Why arent you guys replying?

OpenStudy (wolf1728):

Just taking a guess you could mention derangements - you need to have all the names NOT match their own names.

OpenStudy (adrynicoleb):

I have no idea David. Sorry. ._. Listen to the smart people :P

OpenStudy (davidusa):

thanks for trying adrynicoleb. oh well

OpenStudy (wolf1728):

I can't get a formula - I'm just searching for what topic this might be

OpenStudy (davidusa):

@adrynicoleb can u fan me i want to pm u

OpenStudy (wolf1728):

I'm guesing that the solution of 9 is found by determining the derangements of 4. (A derangement is a sequence in which each member is out of order) 1234 is NOT a derangement 4312 IS

OpenStudy (adrynicoleb):

About math? @DavidUsa

OpenStudy (wolf1728):

From Wikipedia (can you use the link as a answer?) the derangements for 3 people 2 4 people 9 5 people 44 6 people 265 7 people 1854 8 people 14833 9 people 133496 From Wikipedia http://en.wikipedia.org/wiki/Derangement

OpenStudy (davidusa):

Yeah @adrynicoleb

OpenStudy (davidusa):

i dont understand wolfe.

OpenStudy (adrynicoleb):

Well, I suck at math. So that wouldn't work out very well...

OpenStudy (davidusa):

not a question

OpenStudy (wolf1728):

I have found a formula !!! http://math.illinoisstate.edu/day/courses/old/305/contentderangements.html

OpenStudy (davidusa):

wolf can u summarize it here? im in 6th grade thats gonna give me adhd

OpenStudy (davidusa):

all that text

OpenStudy (adrynicoleb):

I am not good at anything related to math...

OpenStudy (davidusa):

not a question i want to talk to u just pm u@adry

OpenStudy (davidusa):

ok adry?

OpenStudy (wolf1728):

WOW they are asking this fro the 6th grade? Derangements??????? The total way the names can be drawn = number of people factorial 4! means factorial = 4*3*2*1 = 24 That's ALL the ways ALL the names are drawn Now to figure the derangements.

OpenStudy (davidusa):

no im advanced math dude. this is mathcounts not classwork. i got 5th individual for chapter.

OpenStudy (wolf1728):

Derangements for 4 people = 4! *[1 - (1/1!) + (1/2!) - (1/3!) + (1/4!)] = 24 * (.5 -.33333333333 + 1/24) = 24 * .375 = 9

OpenStudy (wolf1728):

Made a mistake - 1/3! = 1/6 which equals .1666666666666666 All that's needed is to change in my nswer is to change the .333333333333 to .166666666666666

OpenStudy (davidusa):

i dont understand but i probably wont anyway so u can just leave now.

OpenStudy (davidusa):

@adrynicoleb can u fan me? i want to pm you.

OpenStudy (wolf1728):

Okay - I'm glad I found the formula - send a message if you want it explained some more. Glad I could help you out.

OpenStudy (davidusa):

ok bye wolf